OpenStudy (anonymous):

Express as a single natural log: 4lnx+3lny-5lnz

OpenStudy (zehanz):

You probably know these rules: ln(a)+ln(b)=ln(ab) ln(a)-ln(b)=ln(a/b) k*ln(a)=ln(a^k) Use all three rules here (last one first).

OpenStudy (anonymous):

\[\ln \frac{ x^{4}y^{3} }{ z^{5} }\]

OpenStudy (anonymous):

how does that look?

OpenStudy (zehanz):

Great!

OpenStudy (zehanz):

Just follow the rulez...

OpenStudy (anonymous):

nice! thanks for the assistance on that one! :D

OpenStudy (zehanz):

YW!